# Compressed sensing quantum state tomography for qudits: A comparison of Gell-Mann and Heisenberg-Weyl observable bases

### v1

@@ I Introduction
With the rapid development of quantum processors, quantum circuits, and experimental platforms, understanding the quantum states they generate has become a crucial research topic in quantum information science. A key tool for this purpose is quantum state tomography (QST), which reconstructs the density matrix of an unknown quantum state by measuring expectation values of a complete set of operators. QST has been employed to characterize quantum states in experimental platforms such as trapped ions [1], solid-state qubits [2], and photonic systems [3]. However, QST faces a scalability challenge: the number of required measurement operators grows exponentially with the number of qubits $n$. One approach to address this challenge is compressed sensing quantum state tomography (CS-QST), proposed by D. Gross $et$ ${al}$ [4]. They demonstrated that for a multi-qubit system, the density matrix can be approximately reconstructed using the singular value thresholding (SVT) algorithm [5], requiring only $O(rd\log^{2}d)$ measurement bases instead of the original $d^{2}$, where $r$ and $d$ represent the rank and dimension of the matrix, respectively. QS-QST has since been implemented in photonic [6] and ion-trap [7] systems.
@@
Naturally, QST for qudits requires that the matrix representation of measurement operators has the appropriate dimension; for instance a single qudit with $k=3$, or qutrit, requires a matrix representation of dimension three. The choice of operator basis is not unique and may be motivated by theoretical and experimental considerations. While the standard choice in power-of-two dimension is the Pauli operator basis, two common bases for qudit tomography of arbitrary $k$-dimensional systems are the generalized Gell-Mann (GGM) [23] and Heisenberg-Weyl observable (HWO) [24] bases, which have also been used experimentally in Ref. [25] (GGM) and Ref. [26] (HW). A key distinction between these bases is given by their coherence properties, as defined in Ref. [27]. Coherence properties underlie the mathematical proof of quantum compressed sensing (later refined in terms of the “restricted isometry property”[28, 29]). In the GGM basis, the coherence values grow with $k$, owing to the diagonal operators in particular, whereas in the HWO basis they are independent of $k$ (and all elements share the same spectral norm); consequently, only the latter rigorously meet the hypothesis of the CS-QST theorem in Ref. [27]. Additionally, Asadian $et$ $al.$ [24] noted that all HW observables have maximal rank, which should make them more “universal”, i.e., capable of encoding a certain amount of information characterizing any kind of quantum states In this context, it is worth comparing how different choices of basis affect the efficiency of practical implementations of CS-QST of higher-dimensional density matrices in qudit systems; this ought to be seen as a complementary analysis to the mathematically rigorous results of Ref. [27]. Crucially, providing quantitative insights into the above question can guide future QST experiments, given that setting up the measurement of “theoretically ideal” observables may be extremely costly or outright impossible, and has to be weighed against using an alternative, easier to implement, observable basis.
@@ II.2.1 Generalized Gell-Mann basis
$\footnotesize\begin{split}&\lambda_{1}=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&0\end{pmatrix},\quad\lambda_{2}=\begin{pmatrix}0&-i&0\\ i&0&0\\ 0&0&0\end{pmatrix},\quad\lambda_{3}=\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&0\end{pmatrix}\\ &\lambda_{4}=\begin{pmatrix}0&0&1\\ 0&0&0\\ 1&0&0\end{pmatrix},\quad\lambda_{5}=\begin{pmatrix}0&0&-i\\ 0&0&0\\ i&0&0\end{pmatrix}\\ &\lambda_{6}=\begin{pmatrix}0&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix},\quad\lambda_{7}=\begin{pmatrix}0&0&0\\ 0&0&-i\\ 0&i&0\end{pmatrix},\quad\lambda_{8}=\frac{1}{\sqrt{3}}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&-2\end{pmatrix}\end{split}$ (7)
The appendix V.1 also provides the $k=4$ (quartit) case explicitly.
@@ II.2.2 Heisenberg-Weyl observable basis
$\begin{split}&W(0,0)=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix},\hskip 5.69054ptW(0,1)=\begin{pmatrix}0&\chi^{*}&\chi\\ \chi&0&\chi^{*}\\ \chi^{*}&{\chi}&{0}\end{pmatrix}\\ &W(0,2)=\begin{pmatrix}0&\chi&\chi^{*}\\ \chi^{*}&0&\chi\\ \chi&\chi^{*}&0\end{pmatrix}\\ &W(1,0)=\begin{pmatrix}\chi+\chi^{*}&0&0\\ 0&\chi\theta+\chi^{*}\theta^{*}&0\\ 0&0&\chi\theta^{*}+\chi^{*}\theta\end{pmatrix}\\ &W(1,1)=\begin{pmatrix}0&-\chi^{*}\theta&-\chi\theta\\ -\chi\theta^{*}&0&-\chi^{*}\\ -\chi^{*}\theta^{*}&-\chi&0\end{pmatrix}\\ &W(1,2)=\begin{pmatrix}0&\chi\theta^{*}&\chi^{*}\theta^{*}\\ \chi^{*}\theta&0&\chi\\ \chi\theta&\chi^{*}&0\end{pmatrix}\\ &W(2,0)=\begin{pmatrix}\chi+\chi^{*}&0&0\\ 0&\chi\theta^{*}+\chi^{*}\theta&0\\ 0&0&\chi\theta+\chi^{*}\theta^{*}\end{pmatrix}\\ &W(2,1)=\begin{pmatrix}0&\chi^{*}\theta^{*}&\chi\theta^{*}\\ {\chi}\theta&0&{\chi^{*}}\\ \chi^{*}\theta&\chi&0\end{pmatrix}\\ &W(2,2)=\begin{pmatrix}0&\chi\theta&\chi^{*}\theta\\ \chi^{*}\theta^{*}&0&\chi\\ \chi\theta^{*}&\chi^{*}&0\end{pmatrix},\end{split}$ (12)
@@ II.3 Coherence
Figure 1: The minimum coherence $\nu_{\rm min}$, defined by Eq. (17), is plotted as a function of the dimension $d=k^{N}$, comparing the GGM and HWO bases. Panels (a) and (b) correspond to $N=1$ and $N=2$, respectively, with qudit dimensions ranging from $k=2$ to $k=9$. The insets show magnified views of the HWO results in each panel. Note that for $k=2$, both bases reduce to the Pauli basis.
@@ III.2 Simulation setup and procedure
Starting from the sample matrix $\mathcal{P}_{\Omega}(\bm{M})$, we apply the SVT algorithm [5] to reconstruct the density matrix. A brief review of the SVT algorithm, along with the specific parameters used in our simulations, is provided in Appendix V.3. The reconstructed matrix is denoted as $\bm{X}$, and its accuracy in reproducing the original density matrix is evaluated using two standard metrics, fidelity and trace distance, defined as follows:
$F(\rho,\bm{X})=\frac{{\mathrm{Tr}}(\sqrt{\sqrt{\rho}\bm{X}\sqrt{\rho}})^{2}}{{\mathrm{Tr}(\rho)\mathrm{Tr}(\bm{X})}}{,}$ (25)
$D(\rho,\bm{X})=\frac{1}{2}\mathrm{Tr}|\rho-\bm{X}|{.}$ (26)
@@
We repeat the above procedure 50 times and perform statistical analysis on the results. First, we compute the trace of each reconstructed matrix $\bm{X}$, and only consider those satisfying $20\leq\mathrm{Tr}[\bm{X}]<$ 2 as valid CS-QST outcomes (in practice, in our calculation, invalid outcomes occur very rarely and for very small $m/d^{2}$  [^1] Typically they are less than $4\%$ for the lowest one or two values of $m/d^{2}$ in the simulations presented in Sec. III.3, except for one instance in which they are $8\%$, namely GGM basis, $k=9$ two-qudit states, $m/d^{2}=0.02$. ). The mean values are then computed using these selected samples. Error bars are defined as the range of ±$\pm 1$ standard deviation, calculated as follows:
$\begin{split}&S_{\mathrm{F}}=\sqrt{\frac{1}{{N_{\rm reps}}-1}\sum_{{\mu}=1}^{N_{\rm reps}}({F}(\rho,\bm{X})_{\mu}-{\bar{F}(\rho,\bm{X})})}{,}\\ \\ &S_{\mathrm{D}}=\sqrt{\frac{1}{{N_{\rm reps}}-1}\sum_{{\mu}=1}^{N_{\rm reps}}({D}(\rho,\bm{X})_{\mu}-{\bar{D}(\rho,\bm{X})})}{,}\end{split}$ (27)
where $N_{\rm reps}$ is the number of valid repetitions, and $\bar{F},\bar{D}$ denote the sample means of the fidelity ${F(\rho,\bm{X})}$ and the trace distance ${D(\rho,\bm{X})}$, respectively. The subscript $\mu$ indicates the value obtained in the $\mu$-th repetition. Note that $N_{\rm reps}$ remains nearly 50, as most samples are valid.
@@ III.3 Main results
Figure 2: Fidelity $F(\rho,\bf{X})$ and trace distance $D(\rho,\bm{X})$ between the target two-qudit state $\rho$ and the matrix $\bm{X}$ obtained via CS-QST, for qudit dimensions (a) $k=3$, (b) 6, (c) 9, (d) 12, and (e) 15. The horizontal axis represents the number of randomly selected measurement operators $m$, scaled by the total number of basis elements $d^{2}=k^{4}$.
@@ III.3.1 Two-qudit states
Figure 3: (a) The number of measurements $m_{95\%}$ required to achieve a fidelity of at least 95% in CS-QST of two-qudit states, plotted as a function of the qudit dimension $k$, along with the corresponding fitted curves. The fitted parameters for the GGM basis are $p_{0}=-0.167,p_{1}=20.786,p_{2}=64.676$, and $p_{3}=63.943$; for the HWO basis, $p_{0}=0.027$ and $p_{1}=7.821$. (b) The same data as in (a), but with the vertical axis rescaled by the total number of measurement operators $d^{2}=k^{4}$, showing the relative measurement fraction $m_{95\%}/d^{2}$.
@@
In Fig. 3(a), we observe that when the qudit dimension $k$ is small ($k\lesssim 5$), the number of measurements required to reach the fidelity of at least 95% is comparable for both bases. However, as $k$ increases, a clear discrepancy begins to emerge. The HWO basis consistently requires fewer measurements than the GGM basis to achieve the same level of fidelity, and this difference becomes more pronounced with increasing $k$. These results indicate that the HWO basis offers a more efficient approach to quantum state reconstruction in higher-dimensional qudit systems. This is consistent with expectations based on the coherence structure of the measurement bases, described in Sec. II.3. The HWO basis exhibits nearly constant minimum coherence $\nu_{\rm min}$ regardless of the qudit dimension $k$, whereas the coherence for the GGM basis increases as $(k-1)^{N}$. This increasing coherence makes the GGM basis less favorable in high-dimensional systems, where the HWO basis provides a more stable and efficient reconstruction framework.
@@
The functional form of the required number of measurements $m_{95\%}$ for the GGM and HWO bases, obtained by least-squares fitting, can be expressed as follows:
${m_{95\%}^{\mathrm{GGM}}=k^{2}(p_{0}\log^{2}{(k^{2})}+p_{1})+p_{2}k+p_{3}}$ (28)
with $p_{0}=-0.167,\leavevmode\nobreak\ p_{1}=20.786,\leavevmode\nobreak\ p_{2}=64.676$, and $p_{3}=63.943$;
${m_{95\%}^{\mathrm{HW}}=k^{2}(p_{0}\log^{2}{(k^{2})}+p_{1})}$ (29)
@@
with $p_{0}=0.027$ and $p_{1}=7.821$. Among several candidate fitting models, the above forms were selected based on the Akaike Information Criterion (AIC) [34], which quantitatively evaluates the trade-off between goodness of fit and model complexity (number of fitting parameters). Other candidate functions and their corresponding AIC values are reported in Appendix V.4. These results are broadly consistent with the expected scaling law $O(rd\log^{2}d)$ in compressed sensing [4], as both fitted forms take the structure $k^{2}\left(p_{0}\log^{2}(k^{2})+p_{1}\right)$. However, the fitted values of $p_{1}$ are significantly larger than those of $p_{0}$, indicating that the scaling is closer to $k^{2}$ in practice, at least in the range of dimensions at hand. Moreover, the GGM case includes an additional linear term in $k$, which may reflect the coherence difference between diagonal and off-diagonal elements in the GGM basis, as discussed in Sec.II.3.
We now turn to the analysis of $m_{95\%}/d^{2}$, which provides an important complementary perspective by quantifying the measurement cost relative to the total operator space dimension. Figure 3(b) shows how this relative measurement cost varies with the qudit dimension $k$ for both the GGM and HWO bases. The ratio $m_{95\%}/d^{2}$ decreases approximately as $1/k^{2}$ for both bases, as evident from the fitted form of $m_{95\%}$. This trend indicates that the relative number of measurements required becomes smaller as the qudit dimension increases, demonstrating that CS-QST effectively compresses redundant information in high-dimensional operator spaces.

### v2

@@ I Introduction
With the rapid development of quantum processors, quantum circuits, and experimental platforms, understanding the quantum states they generate has become a crucial research topic in quantum information science. A key tool for this purpose is quantum state tomography (QST), which reconstructs the density matrix of an unknown quantum state by measuring expectation values of a complete set of operators. QST has been employed to characterize quantum states in experimental platforms such as trapped ions [1], solid-state qubits [2], and photonic systems [3]. However, QST faces a scalability challenge: the number of required measurement operators grows exponentially with the number of qubits $n$. One approach to address this challenge is compressed sensing quantum state tomography (CS-QST), proposed by D. Gross $et$ ${al}$ [4]. They demonstrated that for a multi-qubit system, the density matrix can be approximately reconstructed using the singular value thresholding (SVT) algorithm [5], requiring only $O(rd\log^{2}d)$ measurement bases instead of the original $d^{2}$, where $r$ and $d$ represent the rank and dimension of the matrix, respectively. CS-QST has since been implemented in photonic [6] and ion-trap [7] systems.
@@
Naturally, QST for qudits requires that the matrix representation of measurement operators has the appropriate dimension; for instance a single qudit with $k=3$, or qutrit, requires a matrix representation of dimension three. The choice of operator basis is not unique and may be motivated by theoretical and experimental considerations. While the standard choice in power-of-two dimension is the Pauli operator basis, two common bases for qudit tomography of arbitrary $k$-dimensional systems are the generalized Gell-Mann (GGM) [23] and Heisenberg-Weyl observable (HWO) [24] bases, which have also been used experimentally in Ref. [25] (GGM) and Ref. [26] (HWO). A key distinction between these bases is given by their coherence properties, as defined in Ref. [27]. Coherence properties underlie the mathematical proof of quantum compressed sensing (later refined in terms of the “restricted isometry property”[28, 29]). In the GGM basis, the coherence values grow with $k$, owing to the diagonal operators in particular, whereas in the HWO basis they are independent of $k$ (and all elements share the same spectral norm); consequently, only the latter rigorously meet the hypothesis of the CS-QST theorem in Ref. [27]. Additionally, Asadian $et$ $al.$ [24] noted that all HW observables have maximal rank, which should make them more “universal”, i.e., capable of encoding a certain amount of information characterizing any kind of quantum states. In this context, it is worth comparing how different choices of basis affect the efficiency of practical implementations of CS-QST of higher-dimensional density matrices in qudit systems; this ought to be seen as a complementary analysis to the mathematically rigorous results of Ref. [27]. Crucially, providing quantitative insights into the above question can guide future QST experiments, given that setting up the measurement of “theoretically ideal” observables may be extremely costly or outright impossible, and has to be weighed against using an alternative, easier to implement, observable basis.
@@ II.2.1 Generalized Gell-Mann basis
$\begin{split}&\lambda_{1}=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&0\end{pmatrix},\quad\lambda_{2}=\begin{pmatrix}0&-i&0\\ i&0&0\\ 0&0&0\end{pmatrix},\\ &\lambda_{3}=\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&0\end{pmatrix},\quad\lambda_{4}=\begin{pmatrix}0&0&1\\ 0&0&0\\ 1&0&0\end{pmatrix},\\ &\lambda_{5}=\begin{pmatrix}0&0&-i\\ 0&0&0\\ i&0&0\end{pmatrix},\quad\lambda_{6}=\begin{pmatrix}0&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix},\\ &\lambda_{7}=\begin{pmatrix}0&0&0\\ 0&0&-i\\ 0&i&0\end{pmatrix},\quad\lambda_{8}=\frac{1}{\sqrt{3}}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&-2\end{pmatrix}.\end{split}$ (7)
Appendix V.1 also provides the $k=4$ (quartit) case explicitly.
@@ II.2.2 Heisenberg-Weyl observable basis
$\begin{split}&W(0,0)=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix},\quad W(0,1)=\begin{pmatrix}0&\chi^{*}&\chi\\ \chi&0&\chi^{*}\\ \chi^{*}&{\chi}&{0}\end{pmatrix},\quad W(0,2)=\begin{pmatrix}0&\chi&\chi^{*}\\ \chi^{*}&0&\chi\\ \chi&\chi^{*}&0\end{pmatrix},\\ &W(1,0)=\begin{pmatrix}\chi+\chi^{*}&0&0\\ 0&\chi\theta+\chi^{*}\theta^{*}&0\\ 0&0&\chi\theta^{*}+\chi^{*}\theta\end{pmatrix},\quad W(1,1)=\begin{pmatrix}0&-\chi^{*}\theta&-\chi\theta\\ -\chi\theta^{*}&0&-\chi^{*}\\ -\chi^{*}\theta^{*}&-\chi&0\end{pmatrix},\quad W(1,2)=\begin{pmatrix}0&\chi\theta^{*}&\chi^{*}\theta^{*}\\ \chi^{*}\theta&0&\chi\\ \chi\theta&\chi^{*}&0\end{pmatrix},\\ &W(2,0)=\begin{pmatrix}\chi+\chi^{*}&0&0\\ 0&\chi\theta^{*}+\chi^{*}\theta&0\\ 0&0&\chi\theta+\chi^{*}\theta^{*}\end{pmatrix},\quad W(2,1)=\begin{pmatrix}0&\chi^{*}\theta^{*}&\chi\theta^{*}\\ {\chi}\theta&0&{\chi^{*}}\\ \chi^{*}\theta&\chi&0\end{pmatrix},\quad W(2,2)=\begin{pmatrix}0&\chi\theta&\chi^{*}\theta\\ \chi^{*}\theta^{*}&0&\chi\\ \chi\theta^{*}&\chi^{*}&0\end{pmatrix},\end{split}$ (12)
@@ II.3 Coherence
Figure 1: The minimum coherence $\nu_{\rm min}$, defined by Eq. (17), is plotted as a function of the dimension $d=k^{N}$, comparing the GGM and HWO bases. Panels (a) and (b) correspond to $N=1$ and $N=2$, respectively, with qudit dimensions ranging from $k=2$ to $k=16$. The insets show magnified views of the HWO results in each panel. Note that for $k=2$, both bases reduce to the Pauli basis.
@@ III.2 Simulation setup and procedure
Starting from the sample matrix $\mathcal{P}_{\Omega}(\bm{M})$, we apply the SVT algorithm [5] to reconstruct the density matrix. A brief review of the SVT algorithm, along with the specific parameters used in our simulations, is provided in Appendix V.3. The output matrix is denoted as $\bm{X}$; to obtain a bona fide density matrix, we additionally operate a trace normalization and arrive at the final reconstructed density matrix $\tilde{\bm{X}}$. Its accuracy in reproducing the original density matrix is evaluated using two standard metrics, fidelity and trace distance, defined as follows:
$F(\rho,{\tilde{\bm{X}}})=\frac{{\mathrm{Tr}}(\sqrt{\sqrt{\rho}{\tilde{\bm{X}}}\sqrt{\rho}})^{2}}{{\mathrm{Tr}(\rho)\mathrm{Tr}({\tilde{\bm{X}}})}}{,}$ (25)
$D(\rho,{\tilde{\bm{X}}})=\frac{1}{2}\mathrm{Tr}|\rho-{\tilde{\bm{X}}}|{.}$ (26)
@@
We repeat the above procedure 50 times and perform statistical analysis on the results. First, we compute the trace of each reconstructed matrix $\tilde{\bm{X}}$, and only consider those satisfying $20\leq\mathrm{Tr}[\bm{X}]<$ 2 as valid CS-QST outcomes (in practice, in our calculation, invalid outcomes occur very rarely and for very small $m/d^{2}$  [^1] Typically they are less than $4\%$ for the lowest one or two values of $m/d^{2}$ in the simulations presented in Sec. III.3, except for one instance in which they are $8\%$, namely GGM basis, $k=9$ two-qudit states, $m/d^{2}=0.02$. ). The mean values are then computed using these selected samples. Error bars are defined as the range of ±$\pm 1$ standard deviation, calculated as follows:
$\begin{split}&S_{\mathrm{F}}=\sqrt{\frac{1}{{N_{\rm reps}}-1}\sum_{{\mu}=1}^{N_{\rm reps}}({F}(\rho,{\tilde{\bm{X}}})_{\mu}-{\bar{F}(\rho,{\tilde{\bm{X}}})})}{,}\\ \\ &S_{\mathrm{D}}=\sqrt{\frac{1}{{N_{\rm reps}}-1}\sum_{{\mu}=1}^{N_{\rm reps}}({D}(\rho,{\tilde{\bm{X}}})_{\mu}-{\bar{D}(\rho,{\tilde{\bm{X}}})})}{,}\end{split}$ (27)
where $N_{\rm reps}$ is the number of valid repetitions, and $\bar{F},\bar{D}$ denote the sample means of the fidelity ${F(\rho,{\tilde{\bm{X}}})}$ and the trace distance ${D(\rho,{\tilde{\bm{X}}})}$, respectively. The subscript $\mu$ indicates the value obtained in the $\mu$-th repetition. Note that $N_{\rm reps}$ remains nearly 50, as most samples are valid.
@@ III.3 Main results
Figure 2: Fidelity $F(\rho,{\tilde{\bm{X}}})$ and trace distance $D(\rho,{\tilde{\bm{X}}})$ between the target two-qudit state $\rho$ and the matrix ${\tilde{\bm{X}}}$ obtained via CS-QST, for qudit dimensions (a) $k=3$, (b) 6, (c) 9, (d) 12, and (e) 15. The horizontal axis represents the number of randomly selected measurement operators $m$, scaled by the total number of basis elements $d^{2}=k^{4}$.
@@ III.3.1 Two-qudit states
Figure 3: (a) The relative measurement fraction $m_{95\%}/d^{2}$, obtained by normalizing the number of measurements $m_{95\%}$ required to achieve a fidelity of at least 95% in CS-QST of two-qudit states by the total number of measurement operators $d^{2}=k^{4}$, is plotted as a function of the qudit dimension $k$, along with the corresponding fitted curves. The fitted parameters for the GGM basis are $p_{0}=-0.117,\leavevmode\nobreak\ p_{1}=18.865,\leavevmode\nobreak\ p_{2}=-53.044$, and $p_{3}=42.449$; for the HWO basis, $p_{0}=-0.123,\leavevmode\nobreak\ p_{1}=13.897,\leavevmode\nobreak\ p_{2}=-31.882$, and $p_{3}=20.894$. (b) The same data as in (a), but with the vertical axis rescaled by $d^{2}$, showing the number of measurements $m_{95\%}$
@@
In Fig. 3(a), we first analyze the relative measurement cost $m_{95\%}/d^{2}$, which provides an important perspective by quantifying the number of measurements required relative to the total operator space dimension. The functional form of the required relative measurement cost $m_{95\%}/d^{2}$ for the GGM and HWO bases, obtained by least-squares fitting, can be expressed as follows:
${m_{95\%}^{\mathrm{GGM}}/d^{2}={[}k^{2}(p_{0}\log^{2}{(k^{2})}+p_{1})+p_{2}k+p_{3}}{]}/d^{2}$ (28)
with $p_{0}=-0.117,\leavevmode\nobreak\ p_{1}=18.865$, $p_{2}=-53.044$, and $p_{3}=42.449$;
${m_{95\%}^{\mathrm{HW}}/d^{2}={[}k^{2}(p_{0}\log^{2}{(k^{2})}+p_{1})+p_{2}k+p_{3}}{]}/d^{2}$ (29)
@@
with $p_{0}=-0.123,\leavevmode\nobreak\ p_{1}=13.897$, $p_{2}=-31.882$, and $p_{3}=20.894$. The above forms were selected among several candidate fitting models based on the Akaike Information Criterion (AIC) [34], which quantitatively evaluates the trade-off between goodness of fit and model complexity (number of fitting parameters). Other candidate functions and their corresponding AIC values are reported in Appendix V.4. These results are broadly consistent with the expected scaling law $m\sim O(rd\log^{2}d)$ in compressed sensing [4], as both fitted forms take the structure ${{[}k^{2}\left(p_{0}\log^{2}(k^{2})+p_{1}\right)+p_{2}k+p_{3}{]}/d^{2}}$. However, the fitted values of $p_{1}$ are significantly larger than those of $p_{0}$, indicating that the scaling is closer to $m_{95\%}/d^{2}\sim k^{2}/d^{2}=1/k^{2}$ in practice, at least in the range of dimensions at hand. Moreover, the coefficients in the GGM fitting function, particularly $p_{1},\leavevmode\nobreak\ p_{2}$, and $p_{3}$ are significantly larger than those in the HWO case. This may be a consequence of the coherence difference between diagonal and off-diagonal elements in the GGM basis, as discussed in Sec.II.3.
We next report the absolute number of measurements $m_{95\%}$ required to reach a fidelity of at least 95% in Figure 3(b). When the qudit dimension $k$ is small ($k\lesssim 5$), the required number of measurements appears comparable between the two bases. However, as $k$ increases, a clear discrepancy begins to emerge: the HWO basis consistently requires fewer measurements than the GGM basis, with the gap widening for larger $k$ In fact, the growth rate of the gap is approximately $k^{2}$, as inferred from the fitted curve. These results indicate that the HWO basis offers a more efficient approach to quantum state reconstruction in higher-dimensional qudit systems. This is consistent with expectations based on the coherence structure of the measurement bases, described in Sec. II.3. The HWO basis exhibits nearly constant minimum coherence $\nu_{\rm min}$ regardless of the qudit dimension $k$, whereas the coherence for the GGM basis increases as $(k-1)^{N}$. This increasing coherence makes the GGM basis less favorable in high-dimensional systems, where the HWO basis provides a more stable and efficient reconstruction framework.